OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 3x + 7y = 77 -2x + 7y = 42

OpenStudy (anonymous):

3x + 7y = 77---(1) -2x + 7y = 42 ---(2) (1)-(2), 3x+7y+2x-7y=77-42 5x=35 x=7 Sub. x=7 into (3), 3(7)+7y=77 7y=56 y=8

OpenStudy (anonymous):

got it?